### v2

# Fast Compression of Deterministic Finite Automata

Abstract
We revisit the popular delayed deterministic finite automaton (D2FA) compression algorithm introduced by Kumar et al. [SIGCOMM 2006] for compressing deterministic finite automata (DFAs) used in intrusion detection systems. This compression scheme exploits similarities in the outgoing sets of transitions among states to achieve strong compression while maintaining high throughput for matching.
The D2FA algorithm and later variants of it, unfortunately, require at least quadratic compression time since they compare all pairs of states to compute an optimal compression. This is too slow and, in some cases, even infeasible for collections of regular expression in modern intrusion detection systems that produce DFAs of millions of states.
Our main result is a simple, general framework for constructing D2FA based on locality-sensitive hashing that constructs an approximation of the optimal D2FA in near-linear time. We apply our approach to the original D2FA compression algorithm and two important variants, and we experimentally evaluate our algorithms on DFAs from widely used modern intrusion detection systems. Overall, our new algorithms compress up to an order of magnitude faster than existing solutions with either no or little loss of compression size. Consequently, our algorithms are significantly more scalable and can handle larger collections of regular expressions than previous solutions.
1 Introduction
Signature-based deep packet inspection is a key component of modern intrusion detection and prevention systems. The basic idea is to maintain a collection of regular expressions, called signatures, that correspond to malicious content, violations of security policies, etc., and then match the collection of signatures against the input. In the typical scenario of high-throughput network traffic, the matching must be fast enough to process the input at the network’s speed.
A natural approach to solve this is to construct a deterministic finite automaton (DFA) of the collection of signatures and then simulate it on the input. The DFA matches each character of the input with a single constant time state transition and requires only a single memory access. Unfortunately, DFAs for the collections of signatures used in modern intrusion detection systems are prohibitively large and not feasible for practical implementation [49, 6]. To overcome this space issue while still maintaining fast matching, significant work has been done on compressing DFAs [8, 14, 23, 46, 3, 5, 27, 19, 20, 4, 40, 43, 33, 35, 29, 9, 28, 7, 30, 32, 21].
In this paper, we revisit the elegant and powerful delayed DFA compression technique introduced by Kumar et al. [27] and applied in many subsequent solutions [33, 35, 29, 9, 28, 7, 30, 32, 31, 21]. The basic observation is that many states in the DFAs for real-world regular expression collections have similar sets of outgoing transitions, and we can take advantage of this to reduce the space significantly. Specifically, if two states $s$ and $s^{\prime}$ share many such transitions, we can replace these in $s$ with a special default transition to the other state $s^{\prime}$.
Kumar et al. [27] proposed an algorithm to compute an optimal set of default transitions to compress any DFA. The key idea is to compute the similarity of all pairs of states, i.e., the number of shared outgoing transitions. We store these in a complete graph, called the space reduction graph (SRG), on the states with edges weighted by similarity. Finally, we compute a maximum spanning tree on the SRG and use each edge in the tree as a default transition leading to a compressed version of the original DFA called the delayed deterministic finite automaton (D2FA).
To implement matching, we traverse the D2FA similar to the Aho-Corasick algorithm for multi-string matching [2]. When we are at a state $s$ and want to match a character $\alpha$, we first inspect the outgoing transitions at $s$ for a match of $\alpha$. If we find a match, we continue to that state and process the next character in the input, and if not, we follow the default transition to state $s^{\prime}$ and repeat the process from $s^{\prime}$ with character $\alpha$.
Compared to a standard DFA solution, Kumar et al. [27] showed that the D2FA dramatically reduces space by more than 90% on real-world collections of regular expressions while still achieving fast matching performance.
Processing a character D2FA during matching may require following multiple default transitions, thus incurring (as the name suggests) a delay. Minimizing the delay is essential in high-throughput applications, and two important variants of D2FAs that address have been proposed. Kumar et al. [27] gave a modified D2FA construction that, given an integer parameter $L$, limits the maximum path of default transitions to $L$. We refer to this as the longest delay variant of the problem. Alternatively, Becchi and Crowley [7, 9] gave a modified D2FA construction that limits the maximum total number of default transitions traversed on any input string $S$ by $|S|$. We refer to this as the matching delay variant.
The main bottleneck in the above algorithms is computing the similarity of all pairs of states to construct the SRG. If the input DFA contains $n$ states, this requires at least $\Omega(n^{2})$ time and space. This is too slow and, in some cases, even infeasible for collections of regular expression in modern intrusion detection systems that produce DFAs of millions of states.
1.1 Contributions
We present a simple, general framework for fast compression of DFAs with default transitions based on locality-sensitive hashing. We apply our approach to general D2FA compression and the longest delay and matching delay variant and experimentally evaluate our algorithms on collections of regular expressions used in the popular Snort [41], Zeek [36], and Suricata [1] intrusion detection systems (see also www.snort.org, zeek.org, and suricata.io). Overall, we obtain new algorithms that compress up to an order of magnitude faster than existing solutions with either no or little loss of compression size. Consequently, our algorithms are significantly more scalable and can handle larger collections of regular expressions than previous solutions.
Technically, our main idea is to use locality-sensitive hashing to identify approximately similar states quickly. We then add edges between these states weighted by their similarity, producing the sparse space reduction graph (SSRG). We then compute the maximum spanning tree on the SSRG and then the D2FA. The SSRG contains significantly fewer edges, leading to a significant improvement in compression time. While the SSRG approximates the SRG by discarding edges, which may lead to worse compression size, we observe that this loss is negligible experimentally. For the longest delay variant, the previous solution by Kumar et al. [27] also uses a costly heuristic to maintain a bounded diameter maximum spanning forest. If we directly apply our sparsification technique, this heuristic dominates the running time, and we do not experimentally observe a significant speed-up in compression time. Instead, we develop an efficient alternative heuristic that first constructs a maximum spanning tree and then cuts edges to achieve the desired maximum spanning forest with bounded diameter. We show that combining our sparsification with the new heuristic leads to improvements in compression time similar to our other variants with little or no loss of compression size.
1.2 Related Work
Substantial work has been done on compression DFAs. For an overview, see surveys [48, 39]. A popular approach is compressing the set of transitions [33, 46, 3, 5, 27, 28, 7, 19, 20, 9, 35, 29, 4, 40, 30, 32, 43]. This approach includes the popular D2FA algorithm we focus on in this paper. Another approach is to compress the alphabet to reduce the size of the transition table [8, 14, 23, 9, 45]. The main idea is that, if some sets of characters (almost) always cause the same transitions throughout the DFA, they can be replaced by a single character [8, 14, 9, 23]. Alternatively, we can also compress the alphabet by replacing infrequent characters with sequences of frequent characters [45]. Finally, we can also compress the set of states as proposed by Becchi and Cambadi [5]. They showed how states with similar sets of outgoing transitions could be merged into one, thus compressing the set of states.
Using locality-sensitive hashing for fast compression of collections of sets has been used widely in many other contexts [17, 34, 16, 26, 37, 44, 24, 47, 10]. Our work naturally extends this work to fast DFA compression.
2 Preliminaries
Graphs
A graph $G=(V,E)$ is a set of nodes $V$ (also called vertices) and a set of edges $E:V\times V$ between nodes. We call the two nodes of an edge its endpoints. If the edges have direction, i.e. $(u,v)\neq(v,u)$, then the graph is directed, otherwise it is undirected. Edges can have an associated weight, in which case the graph is weighted. Edge can also have an associated label, in which case the graph is labeled. We denote an edge from $u$ to $v$ with label $c$ as ${(u,v)}_{c}$ and say $(u,v)$ is $c$-labeled. A path of length $k$ between two nodes $u_{0}$ and $u_{k}$ is a sequence of nodes $u_{0},\ldots,u_{k}$ such that $(u_{i},u_{i+1})\in E$ for $0\leq i<k$. It can also be viewed as the sequence of edges $(u_{0},u_{1}),\ldots,(u_{k-1},u_{k})$. If $u_{0}=u_{k}$ then $p$ is a cycle. A set of nodes where every pair of nodes have a path between them is called a connected component. If a graph has only one connected component it is connected, otherwise it is disconnected.
A tree is a connected graph that contains no cycles. Due to this property, a tree with $n$ nodes has $n-1$ edges. A graph consisting of several trees is a forest. A node with only one incident edge is called a leaf. For a node $v$ in a tree, its radius is the length of the longest path from $v$ to a leaf. The diameter of a tree is the length of the longest path between any two nodes in the tree. A spanning tree is a subgraph containing all the nodes and is a tree. If the graph is weighted, the value of a spanning tree is the sum of the weights of the edges in the tree. A maximum spanning tree (MST) is a spanning tree of maximum value.
Deterministic Finite Automata
A deterministic finite automaton (DFA) is a 5-tuple $D=(Q,\Sigma,\delta,q_{0},A)$ where $Q$ is a set of states, $\Sigma$ is an alphabet, $\delta:Q\times\Sigma\rightarrow Q$ is a transition function, $q_{0}\in Q$ is the initial state and $A\subseteq Q$ is a set of accepting states. We let $n=|Q|$ denote the number of states. A DFA can be thought of as a labeled directed graph where $Q$ is the set of nodes and each transition $\delta(u,c)=v$ is a labeled, directed edge ${(u,v)}_{c}$. See Figure 1 (A) for an example. For simplicity, we assume every state has exactly one labeled transition for each character in the alphabet, i.e., $\delta$ is total, as in previous work.
Given a string $S$ and a path $p$ in $D$ we say that $p$ matches $S$ if the concatenation of the labels of $p$ equals $S$. A path that starts in $q_{0}$ and ends in $A$ is accepting and $D$ accepts a string $S$ if there exists an accepting path that matches $S$. The language of $D$ is the set of strings it accepts.
Locality-Sensitive Hashing
A family of hash functions is locality-sensitive, for some similarity measure, if the probability of two objects hashing to the same value is high (lower-bounded for some parameter) when they are similar (similarity above some threshold) and, conversely, low when they are dissimilar (see e.g. [22] for formal details). There are different families of locality-sensitive hash functions for different distance or similarity measures, with some of the most popular being simhash [15], MinHash [13] and sdhash [42]. For example, the MinHash of a set is the minimum element according to a uniformly random permutation. The probability that two sets $A$ and $B$ hash to the same value is precisely their Jaccard similarity $(|A\cap B|)/(|A\cup B|)$.
3 Delayed Deterministic Finite Automata
A delayed deterministic finite automaton (D2FA) [27] is a deterministic finite automaton that is augmented with unlabeled default transitions. Formally a D2FA is a 6-tuple $D^{2}=(Q,\Sigma,\delta,q_{0},A,F)$. As for DFAs, $Q$ is the set of states, $\Sigma$ is the alphabet, $\delta$ is the transition functions, $q_{0}\in Q$ is the initial state, and $A\subseteq Q$ is the set of accepting states. The final component $F:Q\rightarrow Q$ is the default transition function. Viewed as a graph, default transitions are $\epsilon$-labeled directed edges, where $\epsilon$ is the empty string, and each state has at most one outgoing default transition. See Figure 1 (C) for an example. To transition from a state $u$ according to a character $c$, we follow a $c$-labeled transition if it exists or otherwise follow the default transition:
$\delta(u,c)=\begin{cases}v&\text{ if }(u,v)_{c}\in D^{2}\\ \delta(F(u),c)&\text{ otherwise }\end{cases}$
Note that for $\delta$ to be well-defined, it must always be possible to reach a state from $u$ with a $c$-labeled transition. This implies that any cycle of default transitions must have an outgoing $c$-labeled transition for any character $c$. To transition from a state $u$ according a string $S=c_{1}\ldots c_{m}$ we recursively transition according to each character:
$\delta(u,c_{1}c_{2}\ldots c_{m})=\delta(\delta(u,c_{1}),c_{2}\ldots c_{m}).$
Given a character $c$ and a path $p=u_{1},\ldots,u_{k}$ we say $p$ matches $c$ if $\delta(u_{1},c)=u_{k}$ and all but the last transition is default, i.e., $F(u_{i})=u_{i+1}$ for $1\leq i<k$. Note that the concatenation of the labels of $p$ equals $c$. Given a string $S=c_{1}\ldots c_{m}$ and a path $p$ we say $p$ matches $S$ if $p$ is the concatenation of the paths matching the individual characters $c_{1},\ldots,c_{m}$. Note that the concatenation of the labels of $p$ is $S$. We define acceptance as before. Two D2FAs are equivalent if they have the same language, and two transitions are equivalent if they have the same destination and label.
Figure 1: Example from [27]. (A) DFA $D$ for regular expression .*((ab+c+)|(cd+)|(bd+e)). Edges to $q_{0}$ are omitted. (B) Space reduction graph for $D$ with edges annotated with similarity. Edges with similarity less than $4$ omitted, except those connecting $q_{2}$ to avoid disconnecting the graph. (C) D2FA equivalent to $D$. All transitions are shown, default transitions are dashed.
[…]
Following a default transition does not consume an input character, which introduces a delay when matching. We define the longest delay of $D^{2}$ to be maximum number of default transitions in any path matching a single character, i.e., the longest delay is the maximum number of default transitions in $D^{2}$ to match any single character. Given a string $S$, we define the matching delay of $S$ in $D^{2}$ to be the number of default transitions in the path starting in $q_{0}$ and matching $S$.
4 Compressing DFAs with Default Transitions
We now review the algorithm of Kumar et al. [27] that compress a DFA $D$ into an equivalent D2FA. Let $D^{2}$ be an initially empty D2FA with the same set of states as $D$. We proceed as follows.
Step 1: Space Reduction Graph
[…]
Step 2: Maximum Spanning Tree
Build a maximum spanning tree over the SRG. Root the spanning tree in a central node, i.e., a node of minimal radius, and direct all edges towards the root to obtain a directed spanning tree.
Step 3: Transitions
For each edge $(u,v)$ in the tree insert the default transition $(u,v)$ into $D^{2}$. Copy every labeled transition from $D$ into $D^{2}$ that is not redundant in $D^{2}$.
Since the SRG weighs the edges by similarity, computing the maximum spanning tree maximizes the overall compression.
Step 1 uses $O(n^{2}|\Sigma|$ to compare all states and construct the SRG. Step 2 constructs the maximum spanning tree with Kruskal’s algorithm [25] in $O(n^{2}\log n)$ time, and finally step 3 takes $O(n|\Sigma|)$ time. In total, the running time is $O(n^{2}\log n+n^{2}|\Sigma|))$ [^1] The running time is not explicitly stated in the paper, but follows from the description..
5 Fast Compression
We now show how to speed up the DFA compression algorithm using locality-sensitive hashing to sparsify the SRG construction in step 1.
Let $D$ be the input DFA and let $r$ and $k$ be two constant, positive integer parameters. We initialize an undirected graph $G=(Q,E)$ where $Q$ is the set of states in $D$ and $E=\{(q_{0},u)\ |\ u\neq v_{0}\}$, i.e., we have a edge between the initial state and all other states. We then add edges to the graph in $r$ rounds, where each round proceeds as follows:
First, pick $k$ unique random characters $c_{1},\ldots,c_{k}\in\Sigma$. Then, for each state $v\in Q$ we construct the sequence of $k$ states $V=\delta(v,c_{1}),\ldots,\delta(v,c_{k})$. We hash $V$ into a single hash value $h(v)$ using a standard hashing scheme of Black et al. [11]. We insert $v$ into a table with key $h(v)$. For each unique hash value $h_{i}$, consider the set of states $C_{i}$ that hash to $h_{i}$. For each state $u\in C_{i}$, we pick another state $v\in C_{i}$ uniformly at random and insert $(u,v)$ into $E$ if it does not already exist.
After $r$ rounds, the algorithm terminates, and we assign weights to each edge of $G$ equal to the similarity of the endpoint states. The resulting graph $G$ is the sparse space reduction graph (SSRG).
The above scheme is inspired the locality-sensitive scheme by Har-Peled et al. [22] for bitstrings w.r.t. Hamming distance. Their solution sample positions from the input bitstrings. In our solution, the sampled positions correspond to the sampled characters from $\Sigma$.
Hashing a state takes $O(k)$ time, and sampling an edge takes constant time. Thus, a round takes $O(kn)$ time and we use $O(rkn)$ time for all rounds. Each round inserts at most $n$ edges and hence the SSRG $G$ has $O(rn)$ edges at the end. We compute the similarity between two states in $O(|\Sigma|)$ time, and hence the final algorithm uses $O(rkn+rn|\Sigma|)=O(n|\Sigma|)$ time.
We plug in the modified Step 1 in the algorithm from Section 4. Step 1 takes $O(n|\Sigma|)$ time and since $G$ has $O(n)$ edges, Step 2 takes $O(n\log n)$ time. Step 3 takes $O(n|\Sigma|)$ time as before. In total, we use $O(n\log n+n|\Sigma|)$ time.
6 Compression with Bounded Longest Delay
We now consider the bounded longest delay variant of D2FA, that is, given a DFA $D$ and an integer parameter $L$, construct a D2FA $D^{2}$ equivalent to $D$, such that the longest delay of $D^{2}$ is at most $L$. Recall that the longest delay of a D2FA is the length of the longest path of default transitions in $D^{2}$.
6.1 Bounded Longest Delay by Constructing Small Trees
We first review the algorithm by Kumar et al. [27]. The algorithm is based on a simple modification to the maximum spanning tree construction in Step 2 of the algorithm from Section 4.
Let $L$ be a parameter. We modify Step 2 by constructing a maximum spanning forest with the constraint that each tree in the forest has a diameter of at most $\Delta=2L$. To do so, we run Kruskal’s algorithm but simply ignore any edges that would cause a tree diameter to exceed $\Delta$. Also, among the edges with maximum similarity, we select one that causes a minimum increase to any tree diameter. After constructing the forest, we root each tree in a central node and direct edges toward each root.
Since each tree has diameter at most $\Delta$ and is rooted in a central node, the final D2FA has a longest delay of at most $\lceil\Delta/2\rceil=L$. To implement the modified Step 2, Kumar et al. [27] maintains the radius of the tree for each node during the maximum spanning forest construction. When we add an edge, we need to merge two trees and potentially update the radius for each node in the resulting tree. Hence, we may need to update $\Omega(n^{2})$ radii in total during the maximum spanning forest construction.
We note that new Step 2 uses $\Omega(n^{2})$ time, whether or not we run it on a sparse or a dense space reduction graph. Hence, we cannot directly apply our sparsification technique from Section 5 to obtain an efficient algorithm. We present a new algorithm in the next section that efficiently combines with sparsification.
6.2 Fast Compression with Bounded Longest Delay
We now present a new algorithm that uses sparsification to efficiently construct D2FAs with bounded longest delay. The idea is to construct a single large maximum spanning tree and then cut edges until each tree in the resulting forest has small diameter.
Given a DFA $D$ and an integer parameter $L$, we proceed as follows.
Step 1: Construct SSRG
Construct a sparse space reduction graph $G$ for $D$ as in Section 5.
Step 2: Construct MST
[…]
Step 3: Cut Edges
Cut a minimum number of edges in $T$ to obtain a forest with each tree diameter at most $\Delta=2L$. We do so using the algorithm of Farley et al. [18] that cuts the necessary edges in a bottom-up traversal of $T$. Then, direct the edges of each tree towards the root.
Step 4: Construct Transitions
As in Step 3 in Section 4, create default transitions along edges in the trees and then copy in every labeled transition from $D$ that is not redundant in $D^{2}$.
After cutting, each tree has a diameter of at most $\Delta=2L$. Since we direct edges toward each root, the longest delay is at most $\lceil\Delta/2\rceil=L$. Step 1 and 4 uses $O(n|\Sigma|)$ time as before. Step 2 uses $O(n\log n)$ time, and Step 3 uses $O(n)$ time. In total, we use $O(n\log n+n|\Sigma|)$ time.
[…]
We found the simple heuristic of the modified weight $w^{\prime}$ performed well in practice. A similar idea was used by Kumar et al [27] in their solution.
Because $T$ is not a maximum spanning tree w.r.t. similarity, the choice of initial node $v_{0}$ affects the total similarity of the final tree. We found that picking $v_{0}$ to be a central node in a MST w.r.t. similarity ($T_{0}$) yielded the best compression in practice.
Note that we cut the minimum number of edges to uphold the diameter constraint. Alternatively, we could cut edges of minimum total similarity, which might result in better compression. However, our approach is simple and fast in practice, and because each edge in the SRG has near-maximum weight, the difference in compression is negligible.
7 Compression with Bounded Matching Delay
We now consider the bounded matching delay variant of D2FA, that is, given a DFA $D$, construct a D2FA $D^{2}$ equivalent to $D$, such that the matching delay of $D^{2}$ is at most $|S|$ on any input string $S$.
7.1 Bounded Matching Delay by the A-DFA Algorithm
We first review A-DFA algorithm by Becchi and Crowley [7, 9]. Let $D$ be an input DFA. For a state $v\in Q$, define the depth of $v$, denoted $d(v)$, to be the length of the shortest path from the initial state $q_{0}$ to $v$. The key idea is only to add default transitions from state $v$ to state $u$ if $d(u)<d(v)$. This implies that the matching delay is at most $|S|$ on any input string $S$ (see, e.g., Aho and Corasick [2]).
Initialize a D2FA $D^{2}$ no default transitions. We proceed as follows.
Step 1: Calculate Depth
Calculate the depth $d(v)$ of each state $v\in Q$ by a breadth-first traversal of $D$.
Step 2: Construct Default Transitions
[…]
Step 3: Construct Labeled Transitions
Copy in every labeled transition from $D$ that is not redundant in $D^{2}$.
Step 1 and 3 uses $O(n|\Sigma|)$ time to traverse $D$. Step 2 uses $O(n^{2}|\Sigma|)$ to compute the similarity of each pair of states. In total, we use $O(n^{2}|\Sigma|)$ time.
7.2 Fast Compression with Bounded Matching Delay
We now speed up the A-DFA algorithm using sparsification. Let $D$ be the input DFA and let $r$ and $k$ be two constant, positive integer parameters. We initialize a D2FA $D^{2}$ with no default transitions, i.e., we set $F(u)=u$ for each $u\in Q$. The algorithm runs in $r$ rounds, where each round proceeds as follows.
[…]
Hashing a state and computing the similarity of the potential new default transition uses $O(k|\Sigma|)$ time. Hence, the full algorithm uses $O(rkn+rn|\Sigma|)=O(n\Sigma)$ time.
8 Experimental Evaluation
We implemented our methods described in the previous section and measured their performance on regular expressions extracted from widely used intrusion detection systems.
8.1 Datasets
We extracted our datasets from regular expressions used in the popular Snort [41], Zeek(formerly Bro) [36], and Suricata [1] intrusion detection systems (see current homepages for these systems at www.snort.org, zeek.org, and suricata.io). The Snort and Zeek datasets are extracted from current versions of the datasets used in most of the previous work.
For each dataset, we extracted prefixes of the rules to generate DFAs of different sizes to explore the scalability of our algorithms on DFAs that have between roughly $1$k to $1$M states. See Table 1 in the appendix for details. As in previous work, we have filtered out some rules that used advanced features. All regular expressions use an ASCII alphabet size of $256$.
8.2 Algorithms Tested
We evaluate the following algorithms.
D2FA.
The algorithm of [27], described in Section 4.
D2FA-Ld.
The algorithm of [27] for bounded longest delay, described in Section 6.1 with parameter $L=2$.
D2FA-Ld-Cut.
The algorithm presented in Section 6.2 for bounded longest delay using the SRG instead of the SSRG.
D2FA-Ld-Cut.
The algorithm presented in Section 6.2 for bounded longest delay using the SRG instead of the SSRG.
D2FA-Md.
The algorithm of [7, 9] for bounded matching delay, described in Section 7.
Sparse-D2FA.
The algorithm presented in Section 5.
Sparse-D2FA-Ld.
As D2FA-Ld using the SSRG instead of the SRG.
Sparse-D2FA-Ld-Cut.
The algorithm presented in Section 6.2 for bounded longest delay.
Sparse-D2FA-Md.
The algorithm presented in Section 7.2 for bounded matching delay.
We use the locality-sensitive hashing scheme from Section 5 without replacement and parameters $k=8$ and $r=512$. We evaluated several locality-sensitive hashing schemes, including the one in Section 5 with replacement and minhash [12] over the set of outgoing transitions with one or $k$ random permutations of the universe. We also evaluated several combinations of $r$ and $k$ and found that $k=8$ gave the best compression size. Increasing $r$ results in better compression size but increases the compression time linearly. Our chosen variant achieved the best combination of compression size and compression time.
In our longest delay variant, we report results for parameter $L=2$. We have experimented with other values of $L$ but did not observe significant differences in the relative performances of the algorithm for this variant. We note that the bounded matching delay variant also appears in a more general version, where we can tune the overhead of the default transitions (see the journal version of the result [9]). The version tested here is the simplest version and leads to the best compression size.
8.3 Setup
Experiments were run on a machine with an Intel Xeon Gold 6226R 2.9GHz processor and $128$GB of memory. The operating system was Scientific Linux 7.9 kernel version 3.10.0-1160.80.1.el7.x86_64. Source code was compiled with g++ version 9.4 with options -Wall -O4. The input to each algorithm is a DFA constructed from a set of regular expressions. We measured the time for constructing an equivalent D2FA for the input DFA, using the clock function of the C standard library.
8.4 Results
We compare the algorithms across the datasets and measure compression time and compression size (number of states in D2FA as a percent of the number of states in the input DFA) for each of the variants (general compression, longest delay, and bounded matching delay). The relative performance of our algorithms is similar across the datasets, so we focus on the results for the Snort dataset shown in Figure 2. The corresponding results for the Suricata and Zeek datasets are in Figures 3 and 4. We point out whenever there are significant differences between observed results across datasets. We note that the absolute compression size varies significantly across the datasets and variants. Most instances are highly compressible to around 10% of the original DFA (many even in the low single-digit percentages). At the same time, a single one (bounded matching delay on the Zeek dataset) compresses to around $50$ percent on the largest DFAs.
Figure 2: Results for the Snort dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
Figure 3: Results for the Suricata dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
Figure 4: Results for the Zeek dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
General Compression
We compare the D2FA and Sparse-D2FA general compression algorithms. We observe that Sparse-D2FA compresses up to an order of magnitude faster than D2FA with either no or little loss of compression size. For DFAs with around $1$k states, the compression time is comparable and increases gradually to roughly an order of magnitude for the largest DFAs.
Compression with Bounded Longest Delay
We compare the D2FA-Ld, Sparse-D2FA-Ld, D2FA-Ld-Cut and Sparse-D2FA-Ld-Cut compression algorithms with bounded longest delay. We observe that D2FA-Ld and Sparse-D2FA-Ld achieve similar compression times as expected. The compression size achieved by Sparse-D2FA-Ld is around 10-15% worse in the Snort dataset, roughly comparable in the Zeek dataset, and around 100% worse in the Suricata dataset. We believe that the bounded diameter approach in Kumar et al. [27] is highly sensitive to the greedy choice of edges at each step, leading to the observed differences in the compression size across the different datasets.
More importantly, we observe that Sparse-D2FA-Ld-Cut compresses up to an order of magnitude faster than all other algorithms. Sparse-D2FA-Ld-Cut also achieves a substantially better compression size than D2FA-Ld and Sparse-D2FA-Ld (except for the Zeek dataset, where the compression size is comparable). Compared to the dense version D2FA-Ld-Cut, Sparse-D2FA-Ld-Cut has either no or little loss of compression size.
Compression with Bounded Matching Delay
We compare the compression algorithms D2FA-Md and Sparse-D2FA-Md with bounded matching delay. We observe that Sparse-D2FA-Md compresses up to an order of magnitude faster than D2FA-Md. The compression size varies depending on the dataset. For the Snort dataset, Sparse-D2FA-Md achieves 10-15% worse compression size for DFAs with more than $10$k states, for the Zeek dataset, Sparse-D2FA-Md is comparable, and for the Suricata dataset, Sparse-D2FA-Md achieves 100% worse compression size for DFAs with more than $10$k states. We note that the absolute compression size varies significantly, which may explain this difference in relative compression size.
9 Conclusion
We presented a simple, general framework for constructing D2FA based on locality-sensitive hashing that constructs an approximation of the optimal D2FA in near-linear time and applied the approach to the D2FA compression algorithm and two important variants. On DFAs from widely used modern intrusion detection systems, we achieved compression times of up to an order of magnitude faster than existing solutions with either no or little loss of compression size.
An interesting open problem is to explore if our new framework can be combined with other DFA compression techniques, such as one mentioned in Section 1.2. Substantial work has been done on compression DFAs and we believe it is likely that combinations with our technique will lead to even faster solutions.
Appendix A Input DFA Characteristics
Source
States
Rules
Average length of rules
% using wildcards (*, +, ?)
% using length restrictions ({,k,+})
Snort
991
12
34.3
50.0
33.3
Snort
1691
14
35.2
57.1
42.9
Snort
5667
15
37.3
60.0
46.7
Snort
8467
22
35.3
59.1
40.9
Snort
15902
50
29.3
46.0
30.0
Snort
39244
61
28.5
41.0
24.6
Snort
83055
101
28.0
53.5
15.8
Snort
144413
102
28.3
53.9
16.7
Snort
246547
128
29.9
49.2
23.4
Snort
523737
134
30.0
48.5
25.4
Snort
1069239
175
37.9
55.4
39.4
Suricata
1066
25
101.6
92.0
0.0
Suricata
2691
28
106.2
89.3
0.0
Suricata
6211
30
112.8
90.0
0.0
Suricata
13577
32
117.2
90.6
0.0
Suricata
32014
38
111.1
86.8
2.6
Suricata
75596
49
105.7
85.7
2.0
Suricata
128696
93
103.6
87.1
1.1
Suricata
273904
101
103.0
87.1
1.0
Suricata
462248
180
101.0
89.4
0.6
Suricata
995152
194
99.1
87.1
0.5
Zeek
2338
5
33.2
60.0
40.0
Zeek
4265
13
46.0
61.5
15.4
Zeek
10137
14
46.2
64.3
14.3
Zeek
12747
17
40.7
52.9
11.8
Zeek
30131
18
42.3
55.6
11.1
Zeek
67933
19
43.7
57.9
10.5
Zeek
127839
20
43.6
60.0
10.0
Zeek
340279
21
44.6
61.9
9.5
Zeek
384949
27
39.6
55.6
7.4
Zeek
1036954
29
39.6
58.6
10.3
Table 1: DFAs used for testing and characteristics of the corresponding set of regular expressions.

### v4

@@
-Fast Compression of Deterministic Finite Automata
+Philip Bille
 Abstract
@@
 Processing a character D2FA during matching may require following multiple default transitions, thus incurring (as the name suggests) a delay.
-Minimizing the delay is essential in high-throughput applications, and two important variants of D2FAs that address have been proposed.
+Minimizing the delay is essential in high-throughput applications, and two important variants of D2FAs that address this problem have been proposed.
 Kumar et al.
@@
 We call the two nodes of an edge its endpoints.
-If the edges have direction, i.e.
-$(u,v)\neq(v,u)$, then the graph is directed, otherwise it is undirected.
+If the edges have direction, i.e., $(u,v)\neq(v,u)$, then the graph is directed, otherwise it is undirected.
 Edges can have an associated weight, in which case the graph is weighted.
-Edge can also have an associated label, in which case the graph is labeled.
+An edge can also have an associated label, in which case the graph is labeled.
 We denote an edge from $u$ to $v$ with label $c$ as ${(u,v)}_{c}$ and say $(u,v)$ is $c$-labeled.
@@
 If $u_{0}=u_{k}$ then $p$ is a cycle.
-A set of nodes where every pair of nodes have a path between them is called a connected component.
-If a graph has only one connected component it is connected, otherwise it is disconnected.
+A set of nodes where every pair of nodes has a path between them is called a connected component.
+If a graph has only one connected component, it is connected, otherwise it is disconnected.
 A tree is a connected graph that contains no cycles.
@@
 To transition from a state $u$ according to a character $c$, we follow a $c$-labeled transition if it exists or otherwise follow the default transition:
-$\delta(u,c)=\begin{cases}v&\text{ if }(u,v)_{c}\in D^{2}\\ \delta(F(u),c)&\text{ otherwise }\end{cases}$
+$\delta(u,c)=\begin{cases}v&\text{ if }(u,v)_{c}\text{ is a edge.}\\ \delta(F(u),c)&\text{ otherwise }\end{cases}$
 Note that for $\delta$ to be well-defined, it must always be possible to reach a state from $u$ with a $c$-labeled transition.
@@
 Note that the concatenation of the labels of $p$ equals $c$.
-Given a string $S=c_{1}\ldots c_{m}$ and a path $p$ we say $p$ matches $S$ if $p$ is the concatenation of the paths matching the individual characters $c_{1},\ldots,c_{m}$.
+Given a string $S=c_{1}\ldots c_{m}$ and a path $p$ we say $p$ matches $S$ if $p$ is the concatenation of paths matching the individual characters $c_{1},\ldots,c_{m}$.
 Note that the concatenation of the labels of $p$ is $S$.
@@
 All transitions are shown, default transitions are dashed.
-[…]
+Given a DFA, we can compress it by replacing sets of equivalent transitions with single default transitions to obtain an equivalent D2FA with fewer total transitions.
+We define the similarity of two states $u$ and $v$, denoted $\textit{sim}(u,v)$, to be their number of equivalent transitions, that is, $\textit{sim}(u,v)=|\{c\in\Sigma\mid\delta(u,c)=\delta(v,c)\}|$.
+See Figure 1 (B) for an illustration.
+Inserting a default transition $(u,v)$ and removing the equivalent transitions from $u$ does not affect the language but saves $\textit{sim}(u,v)-1$ transitions.
+Each transition we can remove without affecting the language we say is redundant.
 Following a default transition does not consume an input character, which introduces a delay when matching.
@@
 Step 1: Space Reduction Graph
-[…]
+Construct the complete, undirected graph over the states of $D$, and to each edge $(u,v)$ assign weight $\textit{sim}(u,v)$. This is the space reduction graph (SRG). See Figure 1 (B) for an example.
 Step 2: Maximum Spanning Tree
@@
 Since the SRG weighs the edges by similarity, computing the maximum spanning tree maximizes the overall compression.
-Step 1 uses $O(n^{2}|\Sigma|$ to compare all states and construct the SRG.
+Step 1 uses $O(n^{2}|\Sigma|)$ to compare all states and construct the SRG.
 Step 2 constructs the maximum spanning tree with Kruskal’s algorithm [27] in $O(n^{2}\log n)$ time, and finally step 3 takes $O(n|\Sigma|)$ time.
@@
 Let $D$ be the input DFA and let $r$ and $k$ be two constant, positive integer parameters.
-We initialize an undirected graph $G=(Q,E)$ where $Q$ is the set of states in $D$ and $E=\{(q_{0},u)\ |\ u\neq v_{0}\}$, i.e., we have a edge between the initial state and all other states.
+We initialize an undirected graph $G=(Q,E)$ where $Q$ is the set of states in $D$ and $E=\{(q_{0},u)\ |\ u\neq q_{0}\}$, i.e., we have a edge between the initial state and all other states.
 We then add edges to the graph in $r$ rounds, where each round proceeds as follows:
@@
 The resulting graph $G$ is the sparse space reduction graph (SSRG).
-The above scheme is inspired the locality-sensitive scheme by Har-Peled et al.
+The above scheme is inspired by the locality-sensitive scheme by Har-Peled et al.
 [23] for bitstrings w.r.t.
 Hamming distance.
-Their solution sample positions from the input bitstrings.
+Their solution samples positions from the input bitstrings.
 In our solution, the sampled positions correspond to the sampled characters from $\Sigma$.
@@
 Step 2: Construct MST
-[…]
+Construct a maximum spanning tree $T_{0}$ over $G$ using Kruskal’s algorithm.
+Pick a central node $v_{0}$ in $T_{0}$ and then discard $T_{0}$.
+Construct a new maximum spanning tree $T$ using Prim’s algorithm [40] with $v_{0}$ as the initial node.
+When we queue a new edge $(u,v)$, where $u$ is the node already in $T$, assign weight $w^{\prime}_{u,v}=\textit{sim}(u,v)-2^{d_{v}}$ where $d_{v}$ is the distance from $v_{0}$ to $v$ in $T$.
 Step 3: Cut Edges
@@
 In total, we use $O(n\log n+n|\Sigma|)$ time.
-[…]
+Each edge $(u,v)$ we cut results in $\textit{sim}(u,v)-1$ more labeled transitions in $D^{2}$, as that default transition is then not constructed.
+Intuitively, the lower the diameter of $T$, the fewer edges we cut to get each tree below the bound.
+Therefore, we use an edge weight that trades similarity for lower diameter, as we observed that the fewer cuts outweighed the lost similarity.
 We found the simple heuristic of the modified weight $w^{\prime}$ performed well in practice. A similar idea was used by Kumar et al [29] in their solution.
@@
 This implies that the matching delay is at most $|S|$ on any input string $S$ (see, e.g., Aho and Corasick [2]).
-Initialize a D2FA $D^{2}$ no default transitions. We proceed as follows.
+Initialize a D2FA $D^{2}$ with no default transitions. We proceed as follows.
 Step 1: Calculate Depth
@@
 Step 2: Construct Default Transitions
-[…]
+For each state $u\in Q$ add default transition $(u,v)$ to $D^{2}$, where $v$ is the state such that $\textit{sim}(u,v)$ is maximum and $d(v)<d(u)$.
 Step 3: Construct Labeled Transitions
@@
 The algorithm runs in $r$ rounds, where each round proceeds as follows.
-[…]
-Hashing a state and computing the similarity of the potential new default transition uses $O(k|\Sigma|)$ time. Hence, the full algorithm uses $O(rkn+rn|\Sigma|)=O(n\Sigma)$ time.
+First, pick $k$ unique random characters $c_{1},\ldots,c_{k}\in\Sigma$.
+Then, for each state $v\in Q$ we construct the sequence $V=\delta(v,c_{1}),\ldots,\delta(v,c_{k})$, and hash $V$ into a single hash value $h(v)$.
+We insert $v$ into a table with key $h(v)$.
+For each unique hash value $h_{i}$, we consider the set of states $C_{i}$ that hash to that value.
+For each state $u\in C_{i}$ we pick another state $v\in C_{i}$ uniformly at random.
+If $v$ has lower depth and the default transition $(u,v)$ compresses better than the current default transition of $u$, i.e., $d(v)<d(u)$ and $\textit{sim}(u,v)>\textit{sim}(u,F(u))$, we update the default transition of $u$ to point to $v$ in $D^{2}$, otherwise, we continue.
+After $r$ rounds, the algorithm terminates and returns the resulting D2FA $D^{2}$.
+Hashing a state and computing the similarity of the potential new default transition uses $O(k|\Sigma|)$ time. Hence, the full algorithm uses $O(rkn+rn|\Sigma|)=O(n|\Sigma|)$ time.
 8 Experimental Evaluation
 We implemented our methods described in the previous section and measured their performance on regular expressions extracted from widely used intrusion detection systems.
+The implementation is available at https://github.com/MaxRishoj/fcomp-dfa.
 8.1 Datasets
@@
 The algorithm of [29] for bounded longest delay, described in Section 6.1 with parameter $L=2$.
-D2FA-Ld-Cut.
-The algorithm presented in Section 6.2 for bounded longest delay using the SRG instead of the SSRG.
 D2FA-Ld-Cut.
@@
 the number of states in the input DFA.
-On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
+On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
 Figure 3: Results for the Suricata dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom).
@@
 the number of states in the input DFA.
-On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
+On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
 Figure 4: Results for the Zeek dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom).
@@
 the number of states in the input DFA.
-On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
+On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
 General Compression
@@
 On DFAs from widely used modern intrusion detection systems, we achieved compression times of up to an order of magnitude faster than existing solutions with either no or little loss of compression size.
-An interesting open problem is to explore if our new framework can be combined with other DFA compression techniques, such as one mentioned in Section 1.2.
-Substantial work has been done on compression DFAs and we believe it is likely that combinations with our technique will lead to even faster solutions.
+An interesting open problem is to explore if our new framework can be combined with other DFA compression techniques, such as the ones mentioned in Section 1.2.
+Substantial work has been done on compression DFAs, and we believe it is likely that combinations with our technique will lead to even faster solutions.
+10 Acknowledgments
+This paper is inspired by earlier work in an MSc.
+thesis [24] supervised by two of the authors.
+We thank the anonymous reviewers of earlier drafts of this article for many valuable comments that improved the quality of the work.
 Appendix A Input DFA Characteristics
-Source
-States
-Rules
-Average length of rules
-% using wildcards (*, +, ?)
-% using length restrictions ({,k,+})
-Snort
-991
-12
-34.3
-50.0
-33.3
-Snort
-1691
-14
-35.2
-57.1
-42.9
-Snort
-5667
-15
-37.3
-60.0
-46.7
-Snort
-8467
-22
-35.3
-59.1
-40.9
-Snort
-15902
-50
-29.3
-46.0
-30.0
-Snort
-39244
-61
-28.5
-41.0
-24.6
-Snort
-83055
-101
-28.0
-53.5
-15.8
-Snort
-144413
-102
-28.3
-53.9
-16.7
-Snort
-246547
-128
-29.9
-49.2
-23.4
-Snort
-523737
-134
-30.0
-48.5
-25.4
-Snort
-1069239
-175
-37.9
-55.4
-39.4
-Suricata
-1066
-25
-101.6
-92.0
-0.0
-Suricata
-2691
-28
-106.2
-89.3
-0.0
-Suricata
-6211
-30
-112.8
-90.0
-0.0
-Suricata
-13577
-32
-117.2
-90.6
-0.0
-Suricata
-32014
-38
-111.1
-86.8
-2.6
-Suricata
-75596
-49
-105.7
-85.7
-2.0
-Suricata
-128696
-93
-103.6
-87.1
-1.1
-Suricata
-273904
-101
-103.0
-87.1
-1.0
-Suricata
-462248
-180
-101.0
-89.4
-0.6
-Suricata
-995152
-194
-99.1
-87.1
-0.5
-Zeek
-2338
-5
-33.2
-60.0
-40.0
-Zeek
-4265
-13
-46.0
-61.5
-15.4
-Zeek
-10137
-14
-46.2
-64.3
-14.3
-Zeek
-12747
-17
-40.7
-52.9
-11.8
-Zeek
-30131
-18
-42.3
-55.6
-11.1
-Zeek
-67933
-19
-43.7
-57.9
-10.5
-Zeek
-127839
-20
-43.6
-60.0
-10.0
-Zeek
-340279
-21
-44.6
-61.9
-9.5
-Zeek
-384949
-27
-39.6
-55.6
-7.4
-Zeek
-1036954
-29
-39.6
-58.6
-10.3
+Source States Rules Average length of rules % using wildcards (*, +, ?) % using length restrictions ({,k,+})
+Snort 991 12 34.3 50.0 33.3
+Snort 1691 14 35.2 57.1 42.9
+Snort 5667 15 37.3 60.0 46.7
+Snort 8467 22 35.3 59.1 40.9
+Snort 15902 50 29.3 46.0 30.0
+Snort 39244 61 28.5 41.0 24.6
+Snort 83055 101 28.0 53.5 15.8
+Snort 144413 102 28.3 53.9 16.7
+Snort 246547 128 29.9 49.2 23.4
+Snort 523737 134 30.0 48.5 25.4
+Snort 1069239 175 37.9 55.4 39.4
+Suricata 1066 25 101.6 92.0 0.0
+Suricata 2691 28 106.2 89.3 0.0
+Suricata 6211 30 112.8 90.0 0.0
+Suricata 13577 32 117.2 90.6 0.0
+Suricata 32014 38 111.1 86.8 2.6
+Suricata 75596 49 105.7 85.7 2.0
+Suricata 128696 93 103.6 87.1 1.1
+Suricata 273904 101 103.0 87.1 1.0
+Suricata 462248 180 101.0 89.4 0.6
+Suricata 995152 194 99.1 87.1 0.5
+Zeek 2338 5 33.2 60.0 40.0
+Zeek 4265 13 46.0 61.5 15.4
+Zeek 10137 14 46.2 64.3 14.3
+Zeek 12747 17 40.7 52.9 11.8
+Zeek 30131 18 42.3 55.6 11.1
+Zeek 67933 19 43.7 57.9 10.5
+Zeek 127839 20 43.6 60.0 10.0
+Zeek 340279 21 44.6 61.9 9.5
+Zeek 384949 27 39.6 55.6 7.4
+Zeek 1036954 29 39.6 58.6 10.3
 Table 1: DFAs used for testing and characteristics of the corresponding set of regular expressions.